The surface area of this cube is 21.66 square yards. What is the value of j?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The surface area of this cube is 21.66 square yards. What is the value of j?
-------
If "j" is an edge, each side has area = j^2
----
The surface area is 6j^2
------
Equation:
6j^2 = 21.66
j^2 = 3.61
j = sqrt(3.61) = 1.9 yrd
-----------
Cheers,
Stan H.
--------------

Answer by Alan3354(69443)

The surface area of this cube is 21.66 square yards. What is the value of j?
------------
j is used in electrical engineering for the sqrt(-1), since i or I is used for current.
----------
Surface area and sq yds is a puzzle, tho.